OpenStudy (anonymous):

Applying the ratio test, what can we say about the convergence of the following series? SEE BELOW

OpenStudy (anonymous):

if \[\frac{ a_{n+1} }{ a _{n} } converges \to p = \frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

and \[b _{n} = 4^{n}a _{n}\]

OpenStudy (anonymous):

I'm suposed to use the ratio test on bn, and give the answer from that, then say if bn converges or diverges